OpenStudy (anonymous):

What is the approximate distance from (6, 5) to (2, 1)? 6 units 5 units 4 units 3 units

OpenStudy (anonymous):

@OrangeMaster

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

I'm still a bit lost @hartnn

hartnn (hartnn):

lets start with x1-x2 = 6-2 = ... ?

OpenStudy (anonymous):

4

hartnn (hartnn):

yes and 4^2 =16 now y1-y2 =5-1 =...

OpenStudy (anonymous):

y1- y2= 5-1

OpenStudy (anonymous):

=4^2

OpenStudy (anonymous):

So 16+ 16= 32

hartnn (hartnn):

yeah, and \(d= \sqrt{16+16} =...?\)

OpenStudy (anonymous):

sqrt of 32= 5.6

hartnn (hartnn):

sqrt 32 approximately is just less than 6 :)

OpenStudy (anonymous):

So a, 6 units

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

While you are here...

OpenStudy (anonymous):

can you help me w one more question?

hartnn (hartnn):

sure

OpenStudy (anonymous):

What is the length of line segment S T if S is at five comma negative one and T is at negative one comma twelve? Round to the nearest hundredth. eleven point seven zero twelve point five three thirteen point six zero fourteen point three two

hartnn (hartnn):

x1-x2 = 5 - (-1) = .... y1-y2 = -1 -12 = ...

OpenStudy (anonymous):

x= 6 y=-13

hartnn (hartnn):

good so \(d =\sqrt {13^2+6^2 }=...\)

OpenStudy (anonymous):

d= 14.31?

OpenStudy (anonymous):

So the answer is D?

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

Also, how do you find the leg of a right triangle if you know one leg and the hypotenuse?

OpenStudy (anonymous):

Thank you sooooooooooo much!!

hartnn (hartnn):

pythagoras theorem \((one ~leg)^2+(other ~leg )^2 = hypotenuse^2\)

OpenStudy (anonymous):

so, hypotenuse^2- one leg^2= other leg^2?

OpenStudy (anonymous):

It's not working.

OpenStudy (anonymous):

Hypotenuse: 10; Leg: 6 would be 4 right?

OpenStudy (anonymous):

Thank you

hartnn (hartnn):

no it would be 8 \(\sqrt{10^2-6^2} = \sqrt{100-36}=8\)

OpenStudy (anonymous):

okay